The length of a rectangle is 8 inches more than twice its width. Its area is 90 in-squared. find the length of the rectangle.
Answer by TimothyLamb(4379)

L = 2w + 8
Lw = 90
---
Lw = 90
(2w + 8)w = 90
2ww + 8w - 90 = 0
---
the above quadratic equation is in standard form, with a=2, b=8 and c=-90
---
to solve the quadratic equation, by using the quadratic formula, copy and paste this:
2 8 -90
into this solver: https://sooeet.com/math/quadratic-equation-solver.php
---
the quadratic has two real roots at:
---
w = 5
w = -9
---
the negative root doesn't fit the problem statement, so use the positive root:
---
w = 5
L = 2w + 8
L = 2*5 + 8
L = 18
---
answer:
L = 18 inches
w = 5 inches
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php